OpenStudy (anonymous):

in summer school -3x+5y=-19 what is x,y slope incperct form

OpenStudy (ivettef365):

y = mx + b where m = slope and b = y-int

OpenStudy (ivettef365):

5y = 3x -19 y = 3/5x -19/5